OpenStudy (anonymous):

I have 2 derivatives that i got wrong on my homework and i want help to solve them. First is find dy/dx if x^2*y + 4*4throot y = xy second is -7t^3 + 3/(t^3) - 3 / (3rdroot t)- And i need to do this without the quotient rule.

OpenStudy (accessdenied):

Do you have the work that you used to get your previous answers?

OpenStudy (anonymous):

\[F(x,y)=x ^{2}y +4\sqrt[4]{y} -xy = 0\] \[dy/dx = - F _{x}(x)/F _{y}\] so find it by implicit diferentiation

OpenStudy (anonymous):

i don't have my steps that i did but i can show you the answers i got. 1st one was (1y --2xy-y ^(1/4))/(x^2 + 4(1/4y)^(-3/4) - x) answer to second one i got -21t^2 + 1/3 (3t)^(-2/3) + 1/3(3t)^(-4/3)

OpenStudy (cwrw238):

x^2*dy/dx + 2xy + 4*1/4 y^(-3/4)*dy/dx = x dy/dx + y solve for dy/dx

OpenStudy (anonymous):

is this revised answer for my first problem correct? (y - 2xy)/(x^2 + y ^(-3/4) - x)

OpenStudy (cwrw238):

x^2*dy/dx +dy/dx* y^9-3/4) - x (dy/dx = y - 2xy yup - thats correct